A solution consisting of acid HA and salt KA If the pH value before adding salt is 3 and after adding salt 5 Calculate the concentration of the added acid


1
Expert's answer
2020-11-12T06:30:47-0500

Your question is not clear and detailed. Well, I presume you meant a substance (salt) was added to an acid with the PH of 3. After adding it the PH of the solution is 5

Change in PH = 5-3

PH of added substance is 2

-log(H+)=2

H+ = 10-2 = 0.01


When next you would send a question, please endeavor to make it clear and detailed. Thank you